OpenStudy (anonymous):

solve using the elimination method. If the system has no solution or an infinite numbers of solutions, state this. 3x - 2y = 11 6x+ 11y = 97

OpenStudy (anonymous):

lol only if it was substitution method... that's the only one i know and still use today.

hero (hero):

Both methods are easy. Substitution is by far the worst method

hero (hero):

3x-2y = 11 6x + 11y = 97 Step 1: Solve both equations for y: y = (11-3x)/-2 y = (97-6x)/11 Set y = y, by elimination and solve for x: 11(11-3x) = -2(97-6x) 121-33x = -194 + 12x 315 = 45x x = 315/45 x = 7 y = 5

OpenStudy (anonymous):

I was wondering how do solve for y in this equation i'm very lost